Twenty metrics of liquidity, Solvency, and Profitability The comparative financial statements of Automotive Solutions Inc. arIncome before income tax $705,000 $916,400 (110,000) Income tax expense (84,200) Net income $806,400 $620,800 AUTOMOTIVE SOLUDecember 31, 2048 and 2017 Dec. 31, 20Y8 Dec. 31, 2017 Assets Current assets: Cash $840,520 $754,570 Temporary investments 1,Bonds payable, 8%, due in 20 years 1,980,000 1,980,000 Total long-term liabilities $3,600,000 $1,980,000 Total liabilities $41. Working capital $ 2. Current ratio 2,595,648 3.2 2.5 3. Quick ratio 4. Accounts receivable turnover 5. Days sales in rece
Answer #1

Answer;

1.Working Capital = Current Assets - Current Liabilities = 3775488 - 1179840 = 2595648

2.Current Ratio = Current Assets / Current Liabilities = 3775488 / 1179840 = 3.2

3.Quick Ratio = Quick Assets / Current Liabilities = ( 840520 + 1272150 + 861400 ) / 1179840 = 2.52

4.Accounts Receivable Turnover = Sales / Avg. Accounts Receivable = 5015100 / ( 861400 + 810300 ) / 2

                                                                                                    = 5015100 / 835850 = 6

5.Days Sales in Receivables = 365 / Accounts Receivable Turnover = 365 / 6 = 60.83 Days

6.Inventory Turnover = Cost of goods sold / Avg. Inventory = 1594320 / ( 642400 + 496400 ) / 2

                                                                                    = 1594320 / 569400 = 2.8

7.Days Sales in Inventory = 365 / Inventory Turnover = 365 / 2.8 = 130.36 Days

8.Debt Ratio = Total Liabilities / Total Assets = 4779840 / 10754640 = 0.44

9.Ratio of Liabilities to Stockholders Equity = 4779840 / 5974800 = 0.80

10.Ratio of Fixed Assets to Long Term Liabilities = 4680000 / 3600000 = 1.3

11.Times Interest earned = (Operating Income + Other Income ) / Interest Expense = (1144180 + 60220 ) / 288000 = 4.18 times

                                     = Operating Income / Interest Expense = 1144180 / 288000 = 3.97 times

12.Times Preferred Dividends Earned = Net Income / Preferred Dividend = 806400 / 13300 = 60.63 times

13.Asset Turnover = Sales / Avg. Total Assets = 5015100 / (10754640 + 8445360 ) / 2 = 5015100 / 9600000 = 0.52

14.Return on Total Assets = Net Income / Avg. Total Assets = 806400 / 9600000 = 8.4%

15.Return on Stockholders Equity = Net Income / Avg. Stockholders Equity = 806400 / ( 5974800 + 5225200 ) / 2

                                                                                                              = 806400 / 5600000 = 14.4%

16.Return on Common Stockholders Equity = (Net Income - Preferred Dividends) / Avg. Common Stockholders Equity

                                                                = ( 806400 - 13300 ) / ( 870000 + 870000 + 4344800 + 3595200 ) / 2

                                                                = 793100 / 4840000 = 16.39%

17.Earning per share = (Net Income - Preferred Dividends) / Outstanding common shares = 793100 / 87000 = 9.12

18.Price Earnings Ratio = Market Price / EPS = 60 / 9.12 = $6.58

19.Dividend per share of Common Stock = Common Dividends / Outstanding common shares = 43500 / 87000 = $0.50

20.Dividend Yield = Dividend per share of Common Stock / Market Price per share = $0.50 / $60 = 0.833%